OpenStudy (tylermckinney16):

sarah borrowed $12,000 to buy a car. If the loan is repaid over 24 months, the monthly payment, including interest, is $550. If the loan is repaid 48 months, the monthly payment, including interest, is $325. How much more would sarah pay if she chooses repaying it over 48 months instead of 24 months.

OpenStudy (tylermckinney16):

@shansha1 Help please!!

OpenStudy (tylermckinney16):

@mangotangochick

OpenStudy (shansha1):

wats sup

OpenStudy (tylermckinney16):

i need help with this question please help.

OpenStudy (shansha1):

idk

OpenStudy (peachpi):

total amount paid with 24 month loan = (24)(550) = total amount paid with 48 month loan = (48)(325) =

OpenStudy (mangotangochick):

First work out 24 months: 24 x 550 = 13200 13200 - 12000 = $1,200 paid extra Then for 48 months: 48 x 325 = 15600 15600 - 12000 = $3,600 paid extra The difference between the two is that 48 month plan is $(3600-1200) more than the 24 month plan.